Balanced equation 2NaOH (aq) + CO2 (g) --> Na2CO3 (aq) + H2O (l)
Calculate the number of grams of carbon dioxide that can be absorbed by complete reaction with a solution that contains 5.00 g of sodium hydroxide.
1
Expert's answer
2012-02-23T09:50:14-0500
Answer: 2.75 g

Explanation:

n(NaOH) = 5g/(40 g*mol-1) = 0.125
moles

n(NaOH):n(CO2) = 2:1 = 0.125:x

x = 0.125/2 = 0.0625
moles

m(CO2) = 0.0625 moles*(44 g*mol-1) = 2.75 g